OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

7(x-1)=5(x+2)+1 .... you have to multiply everything... then add/subtract like turns on same side. turns into 7x-7=5x+11 then you have to do some algebra... turns into 3x=18 turns into x=6

OpenStudy (anonymous):

thank you both :)

OpenStudy (anonymous):

which one is the ans thn 9 or 6?i got 9

OpenStudy (anonymous):

@nafizc3 the answer is 9 your right

OpenStudy (anonymous):

how is it 9 ? just wondering ??

OpenStudy (anonymous):

its solving

OpenStudy (anonymous):

18/3 is 6 ??

OpenStudy (anonymous):

@NikkiC 7x -7= 5x +10+1 2x = 7+10+1 2x = 18 X= 9

OpenStudy (anonymous):

ahh, okay. i see where i messed up now.. lol.. subtraction error... lol

OpenStudy (anonymous):

@NikkiC its okay glad to explain it :)